Zephyr Farming Pty Ltd is considering the purchase of a wind turbine generator in order to generate electricity and to reduce the electricity costs for their offices, which are located in Toowoomba. Currently the business uses 60,000 kilowatt hours (kWh) per quarter (3 months) at an average cost of $0.30 per kwh, supplied by the local coal fired power station. The current required rate of return used to evaluate projects is 6%, with a required payback period of 3 years.
The Queensland government started a scheme to provide an incentive for business to use alternative sources of renewable power. The incentives are 5% immediate reimbursement of the purchase and installation costs. This reimbursement can be paid to the supplier providing and installing the equipment. Therefore, the net cash flow from Zephyr Farming Pty Ltd is the purchase cost plus the installation cost less the 5% incentive back from government to offset these costs.
Project details:
Cost of wind turbine generator          $6,000
Cost to install turbine and generator (by supplier)        $450
Expected cash incentive back from government to offset cost of the panels paid immediately the wind turbine generator installed           5% of total costs
Turbine expected (on average) generated kilowatt              300 hours per month
Generator’s expected life (in years)                   15 years

4. Calculate the net present value (NPV) of the project with a required rate of return at 6%. As the savings are the same each year, use the present value table on the next page for an ordinary annuity. (9 marks)
5. Comment on whether Zephyr should proceed with this project using the Net Present Value method and whether you should base your decision only in this quantitative calculation. (1 mark)
Answer #1

Answer 4:

Initial Investment:

Cost of wind turbine generator = $6,000
Cost to install turbine and generator (by supplier) = $450

Total cost = $6,000 + $450 = $6,450

Expected cash incentive = $6,450 * 5% = $322.50

Net Investment = $6,450 - $322.50 = $6,127.50

Annual cash flow:

Turbine expected (on average) generated kilowatt = 300 hours per month

Annual generated KWH expected = 300 * 12 = 3,600

Annual cash flow = 3,600 * $0.30 = $1,080

PV Factors for a One-Dollar Annuity Discounted at 6% for 15 Periods = [1 - 1/(1 + k) n ] / k = (1 - (1 / (1 + 6%) 15) / 6% = 9.71225

NPV = Annual cash flow * PV factor - Net Initial investment

= $1,080 * 9.71225 - $6127.50

= $4361.73

NPV = $4,361.73

Answer 5:

NPV is positive at $4,361.73 and the project is acceptable.

Zephyr should also consider non quantitative factors while evaluating this project. In general there is a need to look for alternative sources of energy like power generated through wind turbine or solar energy.